Suppose that the market for paint is in equilibrium. The demand for paint is given by Qd=5000-8P. The supply of paint is given by Qs=2000+2P. What is the price elasticity of demand at the market equilibrium?
1
Expert's answer
2015-04-04T11:53:58-0400
Price elasticity of demand (PED) measures the change in the quantity
demanded relative to a change in price for a good or service.
The formula for price elasticity is:
Price Elasticity = (% Change in Quantity) / (% Change in Price)
First of all it is necessary to calculate equilibrium price and demand

Qd = Qs
5000-8P = 2000+2P
10P = 3000
P = 300
Qd = 5000 – 8*300 = 5000 – 2400 = 2600

If price will rise to 400:
% Change in Price = (400 – 300)/300*100% = 33,3%
Qd(new) = 5000 – 8*400 = 5000 – 3200 = 1800
% Change in Quantity = (1800 - 2600)/2600 *100% = -30,7%

Price Elasticity = (% Change in Quantity) / (% Change in Price) =
-30,7%/33,3% = -0,92

Thus, we can say that for every percentage point that prices increase,
the quantity decreases by 0,92 percentage point.